OpenStudy (anonymous):

Write the equation of a circle with a center at 2, -3) and a radius of 1/2 Answers: (x + 2)2 + (y - 3)2 =1/2 (x + 2)2 + (y - 3)2 =1/4 (x - 2)2 + (y + 3)2 =1/4 (x - 2)2 + (y + 3)2 =1/2

OpenStudy (anonymous):

Do you know the formula for a circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What is it

OpenStudy (anonymous):

I believe the answer is d

OpenStudy (anonymous):

(x−h)^2+(y−k)^2=r^2

OpenStudy (anonymous):

plug in the coordinates into the formula

OpenStudy (anonymous):

Don't forget that your radius is also squared

OpenStudy (anonymous):

I did I got D and if the radius is squared it should be D right?

OpenStudy (anonymous):

It would actually be C. Like Shizzle said you have to square your radius. so 1/2^2 = 1/4

OpenStudy (anonymous):

Ohhhhh ok

OpenStudy (anonymous):

C for Close! :) Javi is correct. C for correct, lol

OpenStudy (anonymous):

Thank You